# SUBA-Jet: a new Model for Jets in Heavy Ion Collisions

### v1

@@ Appendix B Sampling the Bethe-Heitler-Gunion-Bertsch seed for $m_{q}=0$
$\Delta\equiv\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ $=$ $B^{+}\times B^{-}$ (59)
@@
$\Delta_{a2}\equiv\Delta(x,\vec{k}_{\perp}=\vec{0},\vec{l}_{\perp}=\vec{0},m_{q})$ $=$ $s(x-x_{\rm min})(x_{\rm max}-x)$ (64)
$=$ $s\left(x-\frac{m_{g}^{2}}{s-(m_{q}+m_{Q})^{2}}\right)\left(1-\frac{(m_{q}+m_{Q})^{2}}{s}-x\right)$
$x_{\rm min}$ $=$ $\frac{m_{g}^{2}}{s-(m_{q}+m_{Q})^{2}}+\mathcal{O}(s^{-2})$ (65)
$x_{\rm max}$ $=$ $1-\frac{(m_{q}+m_{Q})^{2}}{s}+\mathcal{O}(s^{-2})\,.$ (66)
B.1 The cross section and the phase space
For zero mass QGP partons, $m_{q}=0$ and finite gluon mass $m_{g}$, the inelastic Gunion-Bertsch scattering cross section for $Qq\to Qqg$ — with $Q$ being the relativistic projectile quark — can be approximated by model II of ref. Aichelin:2013mra which reads as
@@ B.1 The cross section and the phase space
In the case where both $m_{Q}\neq 0$ and $m_{q}\neq 0$ the full phase space is better expressed in terms of $\vec{k}_{\perp}$, $\vec{l}_{\perp}$ and $\tilde{x}=xP^{+}/P_{\rm tot}^{+}=k^{+}/P_{\rm tot}^{+}$, so that
@@
and $\frac{\mathrm{d}^{5}\sigma_{\rm rad}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}}=\frac{p_{Q}^{+}}{p_{\rm tot}^{+}}\times\frac{\mathrm{d}^{5}\sigma_{\rm rad}}{\mathrm{d}\tilde{x}\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}}$, where the (polarization averaged) squared elastic amplitude $\left|{\cal M}_{\rm el}\right|^{2}$ admits the same expression as in eq. (68), with the general expression of $s_{-}=s-m_{Q}^{2}-m_{q}^{2}$, as in eq. (53). As for the measure, $\Delta$ can be written in an elegant factorized form:
$\Delta\equiv\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})=B^{+}\times B^{-}\,,$ (71)
with expressions of $B^{+}$ and $B^{-}$ given by eqs. (60) and (61). For $m_{q}=0$, one recovers the compact form, eq. (62).
This cross section (70) is both IR and colinear safe thanks to the regulators $\mu$ and $\tilde{m}_{g}$. We work in a generic regime where $s\gg$ these scales. Then, from the $\Theta({\Delta})$ condition, one derives that $l_{\perp}$ should be typically $<\sqrt{s}$ while $k_{\perp}<\sqrt{xs}$. From the respective structures of $\left|\mathcal{M}_{\rm el}\right|^{2}$ and $P_{g}$, one understands that the typical $l_{\perp}\sim\mu$ while the typical $k_{\perp}\sim\tilde{m}_{g}$.
We now provide a precise description of the PS boundaries used later on. As $B^{+}$ is always smaller than $B^{-}$, the original phase space constraint $\Theta(\Delta)$ is identical to $\Theta(B^{+})$ and can be expressed as
@@
which is $\approx l_{y}^{\rm max}$ for small $x$. As $\bar{s}-k_{\perp}^{2}=(1-\tilde{x})\tilde{s}-\frac{k_{\perp}^{2}}{x}$, one sees that both $l_{y}^{\rm max}$ and $l_{x}^{\rm max}$ scale like $\sqrt{s}$ whenever $k_{\perp}\ll\sqrt{x\tilde{s}}$ — a condition that is nearly always satisfied for not too small $x$ — while the $l_{\perp}$ domain shrinks when $k_{\perp}\sim\sqrt{x\tilde{s}}$. The $l_{\perp}$ domain shrinks to 0 when $\bar{s}-{k_{\perp}}^{2}-({m_{q}}+{m_{Q}})^{2}=(1-\tilde{x})\tilde{s}-\frac{k_{\perp}^{2}}{\tilde{x}}-({m_{q}}+{m_{Q}})^{2}=0$, as illustrated on Fig. 41.
Refer to captionRefer to caption
@@
In turn, this equation sets the limits for the $x$-range,
$x_{\rm min}\approx\frac{m_{g}^{2}}{s-(m_{q}+m_{Q})^{2}}\leq x\leq x_{\rm max}\approx 1-\frac{(m_{q}+m_{Q})^{2}}{s}\,.$ (78)
For small $x$, one can thus write
$k_{\perp}^{\rm max}\approx\sqrt{s}\sqrt{x-x_{\rm min}}\,.$ (79)
@@
Now we have obtained the approximate phase space boundaries $x_{\rm min},x_{\rm max}$, $k_{\perp}^{\rm max}(x)$ as well as $l_{x}^{\rm max}(x,k_{\perp})$ and $l_{y}^{\rm max}(x,k_{\perp})$, we will further analyze the impact of approximations performed on the $1/\sqrt{\Delta}$ factor in sampling the cross section. Considering the $1/\sqrt{\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})}$ factor in the cross section, it shows some superficial integrable divergence close to the phase space boundary in the transverse plane. The dependence on $\vec{k}_{\perp}$ and on $\vec{l}_{\perp}$ is difficult to manage when performing the integrals as well as for the Monte Carlo sampling. For these reasons we will approximate $\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ by $\Delta(x,0,0,m_{q})$. Whenever $xs$ is large compared to the typical $l_{\perp}^{2}$ and $k_{\perp}^{2}$ values this is a good approximation because the $k_{\perp}$ and $l_{\perp}$ dependence of $\Delta$ is small and can be neglected compared the dependence of $\left|\mathcal{M}_{\rm el}\right|^{2}\times P_{g}$ on these variables. To be more precise, one can legitimately approximate $\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ by $\Delta(x,0,0,m_{q})$ when in the $\{\vec{l}_{\perp},\vec{k}_{\perp}\}$ space the phase space boundary is located — for a given $x$ — beyond the $\{\vec{l}_{\perp},\vec{k}_{\perp}\}$ values for which the matrix element is maximal, at $l_{\perp}\sim k_{\perp}\approx\mu\approx m_{g}$. Requiring $k_{\perp}^{\rm max}\gtrsim c_{0}\times m_{g}$ is shown from eq. (77) to be equivalent to $x\gtrsim(1+c_{0}^{2})\times x_{\rm min}$, while $l_{\perp}^{\rm max}(k_{\perp}=0)$ increases much faster with $x$ [^25] For $m_{q}=0$ and $k_{\perp}=0$, one has $l_{\perp}^{2}\approx\frac{(\tilde{s}-m_{Q}^{2})^{2}}{4\tilde{s}}$, which is of the order of $m_{g}^{2}$ for $\tilde{s}=2{m_{g}}\sqrt{{m_{g}}^{2}+{m_{Q}}}+2{m_{g}}^{2}+{m_{Q}}$, leading to $x=m_{g}^{2}/(s-m_{Q}^{2}-2m_{g}(m_{g}+\sqrt{m_{g}^{2}+m_{Q}^{2}}))=x_{\rm min}\times\left(1+\frac{2m_{g}(m_{g}+\sqrt{m_{g}^{2}+m_{Q}2})}{s}+\mathcal{O}(s^{-2})\right)$.. Therefore the phase space boundary in the transverse plane will be located beyond the peak region for radiation for $x\gtrsim x_{\rm min}$, legitimating the aforementioned approximation for such $x$-range.
@@
For small $x\sim x_{\rm min}$, the opposite approach is in order. Here the $\{\vec{l}_{\perp},\vec{k}_{\perp}\}$ dependence of $\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ is strong as compared to that of the matrix element $\left|\mathcal{M}_{\rm el}\right|^{2}\times P_{g}$. We then start to analyze this case where the matrix elements can be considered as constant whereas the $\{\vec{l}_{\perp},\vec{k}_{\perp}\}$ dependence of $\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ has to be taken into account explicitly. One notices that $\Delta$ (eq. (59)) can be written as
$\Delta=4x^{2}\bar{s}\times\left[(l_{y}^{\rm max})^{2}-l_{y}^{2}-(1-\frac{k_{\perp}^{2}}{\bar{s}})(l_{x}-\delta)^{2}\right]\,.$ (80)
@@
$\iint\frac{\mathrm{d}l_{x}\,\mathrm{d}l_{y}}{\sqrt{\Delta}}=\frac{\pi}{2x\sqrt{\bar{s}}}\int\mathrm{d}l_{x}\frac{\pi l_{x}^{\rm max}}{x\sqrt{\bar{s}}}\,.$ (81)
@@
which is a factor $4$ smaller then the exact phase space eq. (86). The conclusion is that the approximate phase space $\Phi_{a1}$ or $\Phi_{a2}$ scales like $x-x_{\rm min}$ for all approximations of $\Delta$, but with a smaller slope. [^26] This conclusion stays valid for finite $m_{q}$.
From now on, we will then approximate $\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ by $\Delta_{a2}(x,\vec{0},\vec{0},m_{q})$ – given by eq. (63) – in the expressions (67) and (70) of the radiative cross section. [^27] An intermediate way would be to consider $\Delta(x,\vec{k}_{\perp},\vec{0},m_{q})$ in the measure.. This approximation has no impact when $x\gtrsim x_{\rm min}$ – since the transverse phase space is open enough for $\left|\mathcal{M}_{\rm el}\right|^{2}\times P_{g}$ to vary strongly as compared to $\Delta(x,\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ – and limited impact on the shrinking PS when $x\simeq x_{\rm min}$, as we have just demonstrated. Notice however that the PS condition $\Theta(B^{+})$ is imposed exactly. For the rest of this appendix, we concentrate on the $m_{q}=0$ case. [^28] The general $m_{q}\neq$ case can be treated similarly but implies more tedious expressions. Writing
@@
$\frac{\mathrm{d}^{5}\sigma_{\rm rad}^{sup_{1}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}}\approx\frac{s_{-}}{s}\times\frac{\mathrm{d}\sigma_{\rm el}}{\mathrm{d}^{2}l_{\perp}}\times P_{g}\times{c_{1}(x)}\,.$ (98)
@@
$I_{1}=\frac{C_{A}\alpha_{s}}{\pi}\frac{1-x}{x}\left(\frac{l_{\perp}^{2}+2\tilde{m}_{g}^{2}}{\sqrt{l_{\perp}^{2}(l_{\perp}^{2}+4x^{2}\tilde{m}_{g}^{2})}}\ln\left(\frac{l_{\perp}^{2}\left(\sqrt{1+4x^{2}\tilde{m}_{g}^{2}/l_{\perp}^{2}}+1\right)+4x^{2}\tilde{m}_{g}^{2}}{x^{2}\tilde{m}_{g}^{2}\left(\sqrt{1+4x^{2}\tilde{m}_{g}^{2}/l_{\perp}^{2}}-1\right)}+1\right)-2\right)\,.$ (101)
@@
$\frac{\mathrm{d}^{3}\sigma_{\rm rad}^{{\rm sup}_{2}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}}<\frac{\mathrm{d}^{3}\sigma_{\rm rad}^{{\rm sup}_{3}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}}=\frac{4C_{F}\alpha_{s}^{3}}{\pi}\times c_{2}(x)\frac{1}{(l_{\perp}^{2}+\mu^{2})^{2}}\ln\left(1+\frac{l_{\perp}^{2}}{e\tilde{m}_{g}^{2}}\right)\,,$ (104)
@@
$\frac{\mathrm{d}\sigma_{\rm rad}^{{\rm sup}_{3}}}{\mathrm{d}x}$ $=$ $4C_{F}\alpha_{s}^{3}\times c_{2}(x)\int\mathrm{d}l_{\perp}^{2}\frac{1}{(l_{\perp}^{2}+\mu^{2})^{2}}\ln\left(1+\frac{l_{\perp}^{2}}{e\tilde{m}_{g}^{2}}\right)$ (105)
@@
$\frac{\mathrm{d}\sigma_{\rm rad}^{{\rm sup}_{3}}}{\mathrm{d}x}<\frac{\mathrm{d}\sigma_{\rm rad}^{{\rm sup}_{4}}}{\mathrm{d}x}\coloneqq\frac{4C_{F}\alpha_{s}^{3}}{\mu^{2}}\times\frac{\ln e(\tilde{m}_{g}^{\rm min})^{2}/\mu^{2}}{e(\tilde{m}_{g}^{\rm min})^{2}/\mu^{2}-1}\times c_{2}(x)\,.$ (106)
@@
$\frac{\mathrm{d}\sigma_{\rm rad}^{{\rm sup}_{4}}}{\mathrm{d}x}\approx\frac{4C_{F}\alpha_{s}^{3}}{\mu^{2}}\times\frac{\ln e(\tilde{m}_{g}^{\rm min})^{2}/\mu^{2}}{e(\tilde{m}_{g}^{\rm min})^{2}/\mu^{2}-1}\times\frac{1}{x}\,.$ (107)

### v2

@@ B.1 Definition of the symbols
$\Delta\equiv\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ $=$ $B^{+}\times B^{-}$ (60)
@@
$\Delta_{a2}\equiv\Delta(\tilde{x},\vec{k}_{\perp}=\vec{0},\vec{l}_{\perp}=\vec{0},m_{q})$ $=$ $s^{2}(\tilde{x}-\tilde{x}_{\rm min})(\tilde{x}_{\rm max}-\tilde{x})(\tilde{x}-\tilde{x}^{-}_{\rm min})(\tilde{x}_{\rm max}^{-}-\tilde{x})$ (65)
$\tilde{x}_{\rm min}$ $=$ $\frac{m_{g}^{2}}{s-(m_{q}+m_{Q})^{2}}+\mathcal{O}(s^{-2})$ (66)
$\tilde{x}_{\rm max}$ $=$ $1-\frac{(m_{q}+m_{Q})^{2}}{s}+\mathcal{O}(s^{-2})$ (67)
$\tilde{x}_{\rm min}^{-}$ $=$ $\frac{m_{g}^{2}}{s-(m_{q}-m_{Q})^{2}}+\mathcal{O}(s^{-2})$ (68)
$\tilde{x}_{\rm max}^{-}$ $=$ $1-\frac{(m_{q}-m_{Q})^{2}}{s}+\mathcal{O}(s^{-2})\,.$ (69)
When $m_{q}=0$, we get simplified expressions $\tilde{x}_{\rm min}^{-}=\tilde{x}_{\rm min}^{+}={x}_{\rm min}$ and same for $x_{\rm max}$
$x_{\rm min}$ $=$ $\frac{m_{g}^{2}}{s-m_{Q}^{2}}+\mathcal{O}(s^{-2})$ (70)
$x_{\rm max}$ $=$ $1-\frac{m_{Q}^{2}}{s}+\mathcal{O}(s^{-2})\,.$ (71)
and consequently
$\Delta\equiv\Delta(\tilde{x},\vec{k}_{\perp}=\vec{0},\vec{l}_{\perp}=\vec{0},m_{q}=0)=s^{2}\left(x-\frac{m_{g}^{2}}{s-m_{Q}^{2}}\right)^{2}\left(1-\frac{m_{Q}^{2}}{s}-x\right)^{2}\$ (72)
which agrees up to $\mathcal{O}(s^{-2})$ for $m_{Q}^{2}<<s$ with eq. 65.
B.2 Generic cross section and phase space
For finite mass $m_{Q}$ of the incoming parton and finite mass $m_{g}$ of the radiated gluon but zero mass of the QGP partons, $m_{q}=0$, the inelastic Gunion-Bertsch scattering cross section for $Qq\to Qqg$ — with $Q$ being the relativistic projectile quark — can be approximated by model II of ref. Aichelin:2013mra which reads as
@@ B.2 Generic cross section and phase space
In the case where $m_{q}\neq 0$, the model given by eq. (73) can be extended by expressing the full phase space in terms of $\vec{k}_{\perp}$, $\vec{l}_{\perp}$ and $\tilde{x}=xP_{Q}^{+}/P_{\rm tot}^{+}=k^{+}/P_{\rm tot}^{+}$, so that
@@
and $\frac{\mathrm{d}^{5}\sigma_{\rm rad}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}}=\frac{p_{Q}^{+}}{p_{\rm tot}^{+}}\times\frac{\mathrm{d}^{5}\sigma_{\rm rad}}{\mathrm{d}\tilde{x}\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}}$, where the (polarization averaged) squared elastic amplitude $\left|{\cal M}_{\rm el}\right|^{2}$ admits the same expression as in eq. (74), with the general expression of $s_{-}=s-m_{Q}^{2}-m_{q}^{2}$, as in eq. (54) and where $v_{qQ}=p_{q}\cdot p_{Q}=s_{-}/2$. As for the measure, $\Delta$ can be written in an elegant factorized form:
$\Delta\equiv\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})=B^{+}\times B^{-}\,,$ (77)
with expressions of $B^{+}$ and $B^{-}$ given by eqs. (61) and (62).
This cross section (76) is both IR and colinear safe thanks to the regulators $\mu$ and $\tilde{m}_{g}$. We work in a generic regime where $s\gg$ these scales. Then, from the $\Theta({\Delta})$ condition, one derives that $l_{\perp}$ should be typically $<\sqrt{s}$ while $k_{\perp}<\sqrt{\tilde{x}\tilde{s}}$. From the respective structures of $\left|\mathcal{M}_{\rm el}\right|^{2}$ and $P_{g}$, one understands that the typical $l_{\perp}\sim\mu$ while the typical $k_{\perp}\sim\tilde{m}_{g}$.
B.3 Finite and zero $m_{q}$
B.3.1 Phase space
We now provide a precise description of the phase space boundaries used later on. As $B^{+}$ is always smaller than $B^{-}$, the original phase space constraint $\Theta(\Delta)$ is identical to $\Theta(B^{+})$ and can be expressed as
@@ B.3.1 Phase space
which is $\approx l_{y}^{\rm max}$ for small $\tilde{x}$. As $\bar{s}-k_{\perp}^{2}=(1-\tilde{x})\tilde{s}-\frac{k_{\perp}^{2}}{\tilde{x}}$, one sees that both $l_{y}^{\rm max}$ and $l_{x}^{\rm max}$ scale like $\sqrt{s}$ whenever $k_{\perp}\ll\sqrt{\tilde{x}\tilde{s}}$ — a condition that is nearly always satisfied for not too small $x$ — while the $l_{\perp}$ domain shrinks when $k_{\perp}\sim\sqrt{\tilde{x}\tilde{s}}$. The $l_{\perp}$ domain shrinks to 0 when $\bar{s}-{k_{\perp}}^{2}-({m_{q}}+{m_{Q}})^{2}=(1-\tilde{x})\tilde{s}-\frac{k_{\perp}^{2}}{\tilde{x}}-({m_{q}}+{m_{Q}})^{2}=0$, as illustrated on Fig. 48.
@@
In turn, this equation sets the limits for the $\tilde{x}$-range,
$\tilde{x}_{\rm min}\approx\frac{m_{g}^{2}}{s-(m_{q}+m_{Q})^{2}}\leq\tilde{x}\leq\tilde{x}_{\rm max}\approx 1-\frac{(m_{q}+m_{Q})^{2}}{s}\,.$ (84)
For small $\tilde{x}$, one can thus write
$k_{\perp}^{\rm max}\approx\sqrt{s}\sqrt{\tilde{x}-\tilde{x}_{\rm min}}\,.$ (85)
@@
Now we have obtained the approximate phase space boundaries $\tilde{x}_{\rm min},\tilde{x}_{\rm max}$, $k_{\perp}^{\rm max}(\tilde{x})$ as well as $l_{x}^{\rm max}(\tilde{x},k_{\perp})$ and $l_{y}^{\rm max}(\tilde{x},k_{\perp})$, we will further analyze the impact of approximations performed on the $1/\sqrt{\Delta}$ factor in sampling the cross section. Considering the $1/\sqrt{\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})}$ factor in the cross section, it shows some superficial integrable divergence close to the phase space boundary in the transverse plane. The dependence on $\vec{k}_{\perp}$ and on $\vec{l}_{\perp}$ is difficult to manage when performing the integrals as well as for the Monte Carlo sampling. For these reasons we will approximate $\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ by $\Delta(\tilde{x},0,0,m_{q})$. Whenever $\tilde{x}s$ is large compared to the typical $l_{\perp}^{2}$ and $k_{\perp}^{2}$ values this is a good approximation because the $k_{\perp}$ and $l_{\perp}$ dependence of $\Delta$ is small and can be neglected compared the dependence of $\left|\mathcal{M}_{\rm el}\right|^{2}\times P_{g}$ on these variables. To be more precise, one can legitimately approximate $\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ by $\Delta(\tilde{x},0,0,m_{q})$ when in the $\{\vec{l}_{\perp},\vec{k}_{\perp}\}$ space the phase space boundary is located — for a given $\tilde{x}$ — beyond the $\{\vec{l}_{\perp},\vec{k}_{\perp}\}$ values for which the matrix element is maximal, at $l_{\perp}\sim k_{\perp}\approx\mu\approx m_{g}$. Requiring $k_{\perp}^{\rm max}\gtrsim c_{0}\times m_{g}$ is shown from eq. (83) to be equivalent to claim that $\tilde{x}\gtrsim(1+c_{0}^{2})\times\tilde{x}_{\rm min}$, while $l_{\perp}^{\rm max}(k_{\perp}=0)$ increases much faster with $\tilde{x}$ [^26] For $m_{q}=0$ and $k_{\perp}=0$, one has $l_{\perp}^{2}\approx\frac{(\tilde{s}-m_{Q}^{2})^{2}}{4\tilde{s}}$, which is of the order of $m_{g}^{2}$ for $\tilde{s}=2{m_{g}}\sqrt{{m_{g}}^{2}+{m_{Q}}}+2{m_{g}}^{2}+{m_{Q}}$, leading to $\tilde{x}=m_{g}^{2}/(s-m_{Q}^{2}-2m_{g}(m_{g}+\sqrt{m_{g}^{2}+m_{Q}^{2}}))=\tilde{x}_{\rm min}\times\left(1+\frac{2m_{g}(m_{g}+\sqrt{m_{g}^{2}+m_{Q}2})}{s}+\mathcal{O}(s^{-2})\right)$.. Therefore the phase space boundary in the transverse plane will be located beyond the peak region for radiation for $\tilde{x}\gtrsim\tilde{x}_{\rm min}$, legitimating the aforementioned approximation for such $\tilde{x}$-range.
@@
For small $\tilde{x}\sim\tilde{x}_{\rm min}$, the opposite approach is in order. Here the $\{\vec{l}_{\perp},\vec{k}_{\perp}\}$ dependence of $\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ is strong as compared to that of the matrix element $\left|\mathcal{M}_{\rm el}\right|^{2}\times P_{g}$. We then start to analyze this case where the matrix elements can be considered as constant whereas the $\{\vec{l}_{\perp},\vec{k}_{\perp}\}$ dependence of $\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ has to be taken into account explicitly. One notices that $\Delta$ in eq. (60) can be written as
$\Delta=4\tilde{x}^{2}\bar{s}\times\left[(l_{y}^{\rm max})^{2}-l_{y}^{2}-(1-\frac{k_{\perp}^{2}}{\bar{s}})(l_{x}-\delta)^{2}\right]\,.$ (86)
@@
$\iint\frac{\mathrm{d}l_{x}\,\mathrm{d}l_{y}}{\sqrt{\Delta}}=\frac{\pi}{2\tilde{x}\sqrt{\bar{s}}}\int\mathrm{d}l_{x}=\frac{\pi l_{x}^{\rm max}}{\tilde{x}\sqrt{\bar{s}}}\,.$ (87)
@@
which is a factor $4$ smaller then the exact phase space eq. (92). The conclusion is that the approximate phase space $\Phi_{a1}$ or $\Phi_{a2}$ scales like $x-x_{\rm min}$ for all approximations of $\Delta$, but with a smaller slope. We have checked that this conclusion stays valid for finite $m_{q}$.
From now on, we will then approximate $\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ by $\Delta_{a2}(\tilde{x},\vec{0},\vec{0},m_{q})$ – given by eq. (64) – in the expressions (73) and (76) of the radiative cross section. [^27] An intermediate way would be to consider $\Delta(\tilde{x},\vec{k}_{\perp},\vec{0},m_{q})$ in the measure.. This approximation has no impact when $\tilde{x}\gtrsim\tilde{x}_{\rm min}$ – since the transverse phase space is open enough for $\left|\mathcal{M}_{\rm el}\right|^{2}\times P_{g}$ to vary strongly as compared to $\Delta(\tilde{x},\vec{k}_{\perp},\vec{l}_{\perp},m_{q})$ – and limited impact on the shrinking PS when $\tilde{x}\simeq\tilde{x}_{\rm min}$, as we have just demonstrated. Notice however that the PS condition $\Theta(B^{+})$ is imposed exactly.
B.3.2 Explicit radiative cross section for $m_{q}=0$
We present now the explicit expression for the radiative cross section for $m_{q}=0$, implying $\tilde{x}=x$ [^28] The general $m_{q}\neq 0$ case can be treated similarly but implies more tedious expressions.. Writing
@@ B.3.2 Explicit radiative cross section for $m_{q}=0$
$\frac{\mathrm{d}^{5}\sigma_{\rm rad}^{{\rm sup}_{1}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}\,\mathrm{d}^{2}k_{\perp}}\approx\frac{s_{-}}{s}\times\frac{\mathrm{d}\sigma_{\rm el}}{\mathrm{d}^{2}l_{\perp}}\times P_{g}\times{c_{1}(x)}\,.$ (104)
@@
$I_{1}=\frac{C_{A}\alpha_{s}}{\pi}\frac{1-x}{x}\left(\frac{l_{\perp}^{2}+2\tilde{m}_{g}^{2}}{\sqrt{l_{\perp}^{2}(l_{\perp}^{2}+4\tilde{m}_{g}^{2})}}\ln\left(\frac{l_{\perp}^{2}\left(\sqrt{1+4\tilde{m}_{g}^{2}/l_{\perp}^{2}}+1\right)+4\tilde{m}_{g}^{2}}{\tilde{m}_{g}^{2}\left(\sqrt{1+4\tilde{m}_{g}^{2}/l_{\perp}^{2}}-1\right)}+1\right)-2\right)\,.$ (107)
@@
$\frac{\mathrm{d}^{3}\sigma_{\rm rad}^{{\rm sup}_{2}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}}<\frac{\mathrm{d}^{3}\sigma_{\rm rad}^{{\rm sup}_{3}}}{\mathrm{d}x\,\mathrm{d}^{2}l_{\perp}}=\frac{4C_{F}\alpha_{s}^{3}}{\pi}\times c_{2}(x)\frac{\ln\left(1+\frac{l_{\perp}^{2}}{e\tilde{m}_{g}^{2}}\right)}{(l_{\perp}^{2}+\mu^{2})^{2}}\,,$ (110)
@@
$\frac{\mathrm{d}\sigma_{\rm rad}^{{\rm sup}_{3}}}{\mathrm{d}x}$ $=$ $4C_{F}\alpha_{s}^{3}\times c_{2}(x)\int\mathrm{d}l_{\perp}^{2}\frac{\ln\left(1+\frac{l_{\perp}^{2}}{e\tilde{m}_{g}^{2}}\right)}{(l_{\perp}^{2}+\mu^{2})^{2}}$ (111)
@@
$\frac{\mathrm{d}\sigma_{\rm rad}^{{\rm sup}_{3}}}{\mathrm{d}x}<\frac{\mathrm{d}\sigma_{\rm rad}^{{\rm sup}_{4}}}{\mathrm{d}x}\coloneqq 4C_{F}\alpha_{s}^{3}\times\frac{\ln e(\tilde{m}_{g}^{\rm min})^{2}/\mu^{2}}{e(\tilde{m}_{g}^{\rm min})^{2}-\mu^{2}}\times c_{2}(x)\,.$ (112)
@@
$\frac{\mathrm{d}\sigma_{\rm rad}^{{\rm sup}_{4}}}{\mathrm{d}x}\approx 4C_{F}\alpha_{s}^{3}\times\frac{\ln e(\tilde{m}_{g}^{\rm min})^{2}/\mu^{2}}{e(\tilde{m}_{g}^{\rm min})^{2}-\mu^{2}}\times\frac{1}{x}\,.$ (113)